# Diffusion Reward: Learning Rewards via Conditional Video Diffusion

### v2

@@ Appendix E More Generalization Experiments
Figure 13: More results of reward generalization. Diffusion Reward exhibit better generalization ability than VIPER. Results are means of 5 seeds with standard errors (shaded area).
Appendix F Visualization of Reward and Trajectory in Simulation and Real Robot
Simulation Results. We first visualize the reward curve in 10 simulation tasks in Figure 15. Our proposed reward can greatly distinguish the expert-like and -unlike behaviors. Interestingly, we observe that two door-opening tasks show a return drop at the final execution stage. This may be attributed to the difficulty of modeling the dynamics of the door, suggesting that explicit modeling of environmental dynamics is worth investigating in the future.
Real Robot Results. We collect 20 real robot video trajectories with an Allegro hand, a Franka arm, and a RealSense D435i camera. There is only one task, which is picking up a bowl on the table. 10 of the videos are success trajectories while the other 10 are random trajectories. We train our pipeline on the expert demonstrations and evaluate Diffusion Reward on both expert and random trajectories. Visualization results in Figure 16 show that Diffusion Reward can correctly assign expert demonstrations relatively higher reward and random trajectories lower reward.

### v3

@@ 0.D.3 More Generalization Experiments
Figure 19: More results of reward generalization. Diffusion Reward exhibit better generalization ability than VIPER. Results are means of 5 seeds with 95% standard errors (shaded area).
Figure 20: AMP generalization
Comparison with AMP.
Recall that AMP is excluded from our generalization experiments due to its requirement on target expert data. That said, we design a pseudo-generalization test in Figure 20, where AMP is trained on Door Close and frozen in the Drawer Open. We named it ‘pseudo’ because, though AMP does not require expert videos from the target task (Drawer Open) here, it requires additional online interactions on Door Close tasks, which differs from our default settings. We surprisingly observe a fair performance, possibly because AMP learns to discern the behavior of reaching objects and transfer this ability. Still, our method outperforms others.
Appendix 0.E Visualization of Reward and Trajectory in Simulation and Real Robot
Simulation Results.
We first visualize the reward curve in 10 simulation tasks in Figure 22. Our proposed reward can greatly distinguish the expert-like and -unlike behaviors. Interestingly, we observe that two door-opening tasks show a return drop at the final execution stage. This may be attributed to the difficulty of modeling the dynamics of the door, suggesting that explicit modeling of environmental dynamics is worth investigating in the future.
Real Robot Results.
We collect 20 real robot video trajectories with an Allegro hand, a Franka arm, and a RealSense D435i camera. There is only one task, which is picking up a bowl on the table. 10 of the videos are success trajectories while the other 10 are random trajectories. We train our pipeline on the expert demonstrations and evaluate Diffusion Reward on both expert and random trajectories. Visualization results in Figure 23 show that Diffusion Reward can correctly assign expert demonstrations relatively higher reward and random trajectories lower reward.